Calculate the discharge in litres per second and the time required to discharge 300 000 litres of water from a swimming pool by means of a 76mm diameter short pipe, having a bell shape connection when there is a constant head of water above the centre of the pipe of 2 meters
1
Expert's answer
2019-12-16T11:09:02-0500

The discharge:


"Q=C_d A\sqrt{2gh}"

"Q=(0.97)(0.25\pi (0.076)^2)\sqrt{2(9.81)(2)}=0.02756\frac{m^3}{s}"

"Q=27.56\frac{litre}{s}"

Time:


"T=\frac{V}{Q}=\frac{300000}{27.56}=10885\ s=3.024\ h"